### v2

# Chirally-protected state manipulation by tuning one-dimensional statistics

Abstract
Chiral symmetry is broken by typical interactions in lattice models, but the statistical interactions embodied in the anyon-Hubbard model are an exception. It is an example of a correlated hopping model in which chiral symmetry protects a degenerate zero-energy subspace. Complementary to the traditional approach of anyon braiding in real space, we adiabatically evolve the statistical parameter in the anyon-Hubbard model and we find non-trivial Berry phases and holonomies in this chiral subspace. The corresponding states possess stationary checkerboard pattern in their $N$-particle densities which are preserved under adiabatic manipulation. We give an explicit protocol for how these chirally-protected zero energy states can be prepared, observed, validated, and controlled.
In two dimensions, Abelian braid anyons with fractional exchange statistics arise from the topological analysis of two-body interactions [1, 2, 3, 4, 5, 6], and non-Abelian anyons of various forms have been proposed as the working material for robust topological quantum computing protocols [7, 8]. However, non-standard exchange statistics are not an exclusively two-dimensional phenomenon. Their key features have been proposed and investigated in one-dimensional systems since the beginning of the field [1, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19], leading to experimental proposals [20, 21, 22, 23, 24, 25] and recent realizations in Raman-coupled Bose-Einstein condensates [26, 27]. This, for the first time, opens the possibility to tune the statistical angle of anyons in an experimentally accessible platform.
The anyon-Hubbard model provides the platform for this exploration [15, 28, 29, 30]. It realizes exchange statistics with a statistical angle $\theta$ that interpolates between bosons $\theta=0$ and fermions $\theta=\pi$ are implemented on a one-dimensional lattice using Floquet-manipulated Rb atoms in quantum gas microscopes [31] by density-dependent Peierls phases  [32, 33, 34]. Such phases lead to intriguing effects, including statistically induced phase transitions [15, 28, 35], quasi-condensation at finite momenta [15, 28], emerging Friedel oscillations [29, 36, 30], as well as asymmetrical transport and expansion dynamics. [37, 38].
Although the anyons realized by the anyon-Hubbard model are Abelian, their non-standard exchange statistics reveals topological structures in configuration space which we propose to be exploited for non-Abelian state manipulation. To motivate this, consider that the density-dependent Peierls phases which implement the statistical interaction can induce synthetic magnetic fluxes through plaquettes in configuration space [31, 19]. In the anyon-Hubbard model, these fluxes provide a phase $\exp(\pm i\theta)$ depending on the order in which the particles exchange; see Fig. 1(a). Many models in which underlying canonical particles experience correlated hopping processes share this feature, including [39, 32, 38, 33, 40, 41, 42]. Such correlated hopping processes can be engineered to break parity and time-reversal symmetry [43]. However, correlated hopping models with only nearest-neighbor hopping processes, such as the anyon-Hubbard model, preserve the chiral symmetry associated with bipartite lattices, a symmetry that is broken when non-statistical interactions are included [44, 45]. As a result, theorems about bipartite spatial lattices [46, 47, 48, 49] can be generalized from real space to configuration space in models with chirally symmetric interactions and particle number conservation.
In this article, we show that one-dimensional lattice anyons with only statistical interactions host a degenerate zero energy subspace protected by chiral symmetry. Adiabatically tuning the statistical angle $\theta$ from 00 to $2\pi$ varies all energy levels in the spectrum except for the zero-energy modes, as shown in Fig. 1(b1-b2). In this space, the variation generates nontrivial holonomies [50, 51], i.e.,  unitary transformations similar to braiding non-Abelian two-dimensional anyons around each other [8]. This scheme can be used for chiral-protected non-Abelian state preparation. Building on theorems of Lieb and Sutherland [46, 47], we find that these results hold for the experimentally accessible anyon-Hubbard model and extend to chirally symmetric correlated hopping models of bosons with particle number conservation. In these models, chiral symmetry equips the configuration space with a stationary checkerboard pattern that can be experimentally revealed by the $N$-body density correlations and the spatially off-diagonal correlation functions of zero energy states and the quasi-momentum distribution. We furthermore show how these zero energy states can be prepared from typical initial states and manipulated by steering the statistical angle.
Figure 1: Configuration space representation of the anyon-Hubbard model and its chirally-protected zero energy subspace $\mathcal{H}_{0}$. (a) Bosonic configuration space for $N=2$ particles on $L=6$ sites. Each dot represents a bosonic number state with doubly-occupied sites (black) and singly-occupied (not black) and with positive chirality (black, red hashed) or negative chirality (plain blue). The edges represent hopping processes with amplitude $-J$ (thin), $-\sqrt{2}{J}$ (thick, dotted), and $-\sqrt{2}Je^{i\theta}$ (arrow). Boundary plaquettes have a flux with a density-dependent Peierls phase $\theta$ through them. (b1) Energy spectrum as a function of the statistical angle $\theta$. The zero energy subspace has dimension $d_{0}=3$ for $N=2$ and $L=6$ (red line) without avoided crossings (inset (b2)). (b3) Schematic rotation of $\mathcal{H}_{0}$ by cyclic manipulation of $\theta$. (c) The minimal degeneracy $d_{0}$ of $\mathcal{H}_{0}$.
Chiral symmetry and correlated hopping.— For a particle-number conserving bosonic system, a chiral symmetry corresponds to a unitary operator $\hat{S}$ that anticommutes with the Hamiltonian $[\hat{S},\hat{H}]_{+}=-\hat{H}$ [52, 53, 54]. Such an operator is an involution $\hat{S}^{2}=1$ that partitions the finite-dimensional Hilbert space into chiral eigenspaces $\mathcal{H}=\mathcal{H}_{+}\oplus\mathcal{H}_{-}$ with eigenvalues $\chi=\pm 1$ and dimensions $\dim\mathcal{H}_{\pm}=d_{\pm}$. $\hat{H}$ is anti-block diagonal in the chiral basis and has a symmetric spectrum. Therefore, the zero energy eigenspace $\mathcal{H}_{0}$ has dimension $d_{0}\geq|d_{+}-d_{-}|$ [46, 47, 48, 49, 55]. This minimal degeneracy of the zero energy subspace $d_{0}$ does not depend on the details of the Hamiltonian and is protected against perturbations that preserves chiral symmetry.
For $N$ spinless bosons on a one-dimensional lattice with $L$ sites, the operator $\hat{S}$ that realizes chiral symmetry for nearest-neighbor correlated-hopping models is [44, 45, 53]:
$\hat{S}=\exp\left(i\pi\sum_{k=1}^{L}k\hat{n}_{k}\right),$ (1)
where $\hat{n}_{k}=\hat{b}^{\dagger}_{k}\hat{b}_{k}$ is the number operator on site $k$. The operator $\hat{S}$ transforms bosonic operators as $\hat{S}\hat{b}_{k}\hat{S}=(-1)^{k}\hat{b}_{k}$ and acts like a local gauge transformation in configuration space that assigns opposite chirality to adjacent number states, i.e., number states that differ by a single hop have opposite chirality. See Fig. 1(a) for a depiction of $L=6$ and $N=2$, where the chiral operator partitions the $21$-site configuration space lattice into a checkerboard of sublattices with $d_{+}\!=\!12$ and $d_{-}\!=\!9$ chiral number states, respectively. Chiral symmetry together with particle-number conservation guarantees that this system has at least $d_{0}\!=\!|d_{+}-d_{-}|=3$ zero-energy states with positive chirality, although the specific subspace $\mathcal{H}_{0}\subset\mathcal{H}_{+}$ spanned by these three states depends on the Hamiltonian. Note that additional pairwise zero energy subspace degeneracies can only appear due to other non-Abelian symmetries or special accidental degeneracies. In particular, the non-interacting Bose-Hubbard model with $L=8$ and $N\geq 3$ has additional degeneracies at $E=0$ because of the trigonometric relation $\cos(\pi/9)+\cos(5\pi/9)+\cos(7\pi/9)=0$. These accidental, cyclotomic degeneracies are akin to the Pythagorean degeneracies of the infinite square well. They occur in pairs that split upon variation from $\theta=0$ and are therefore not chirally-protected.
On the zero energy subspace, the chiral operator $\hat{S}$ is promoted to an actual symmetry, because $\mathcal{H}_{0}$ is a subspace of the majority chirality eigenspace $\mathcal{H}_{\pm}$ [46, 47, 49, 55]. This implies that all correlation functions vanish which are not invariant with respect to $\hat{S}$. For example, the operator $\hat{b}^{\dagger}_{j}\hat{b}_{j+1}$ has odd chirality and therefore has a vanishing expectation value for every state in $\mathcal{H}_{0}$. Remarkably, we find for $N<L$ an analogous pattern in the $N$-particle densities. States in the zero energy subspace have support exclusively on the majority sublattice [46, 47, 56], forming an higher-dimensional generalization of a checkerboard pattern. The scenario of spontaneous symmetry breaking in the thermodynamic limit is ruled out by Elitzur’s theorem [57, 58, 59].
For general $N$ and $L$, the dimension of the chirally-protected zero energy subspace can be derived using combinatorics [49, 60, 61] and is depicted in Fig. 1(c):
$d_{0}=\left\{\begin{array}[]{ll}0&\mbox{for $N$ odd and $L$ even}\\ \frac{(\lceil L/2\rceil+\lfloor N/2\rfloor-1)!}{(\lceil L/2\rceil-1)!\lfloor N/2\rfloor!}&\mbox{else,}\end{array}\right.$ (2)
where $\lfloor A\rfloor$ and $\lceil A\rceil$ are the floor and ceiling function of $A$, respectively. The majority chiral subspace is $\mathcal{H}_{+}$ for $N$ even and $\mathcal{H}_{-}$ for $N$ odd and $L$ odd.
Model.— The statistical interactions of the anyon-Hubbard model are chirally symmetric, in contrast to ordinary Hubbard-type interactions. They embody momentum-dependent interactions that are periodic in the statistical angle $\theta$. The model is defined in terms of anyonic operators $\hat{a}_{j}$ with deformed commutation relations that obey fractional exchange statistics
$\hat{a}_{j}\hat{a}_{k}^{\dagger}-e^{-i\theta\mathrm{sgn}(j-k)}\hat{a}^{\dagger}_{k}\hat{a}_{j}$ $=$ $\delta_{jk}$
$\hat{a}_{j}\hat{a}_{k}-e^{i\theta\mathrm{sgn}(j-k)}\hat{a}_{k}\hat{a}_{j}$ $=$ $0$ (3)
The Hamiltonian with $L$ sites takes the form
$H(\theta)=-J\sum_{j=1}^{L-1}\left(\hat{a}_{j+1}^{\dagger}\hat{a}_{j}+\mathrm{h.c.}\right).$ (4)
The model is mapped to canonical bosons by a fractional Jordan-Wigner transformation $\hat{a}_{j}=\hat{b}_{j}e^{i\theta\sum_{l<j}\hat{n}_{l}}$ [15]:
$\hat{H}(\theta)$ $=$ $-J\sum_{j=1}^{L-1}\left(\hat{b}_{j+1}^{\dagger}e^{-i\theta\hat{n}_{j}}\hat{b}_{j}+\mathrm{h.c}\right).$ (5)
This results in a correlated hopping process mediated by a density-dependent Peierls phase.
Properties of the zero energy subspace.— We investigate the properties of the zero energy subspace by exact diagonalization and find the dimension of the zero energy subspace $d_{0}$ typically assumes its minimal value according to Eq. (2) as expected when the only symmetries are Abelian. Additional accidental degeneracies appear for special $N$ and $L$ at $\theta=0$ that are not chirally-protected and unsuitable for adiabatic manipulation.
We first construct a convenient basis for the zero energy subspace at $\theta=0$ (non-interacting bosons) and use this as a experimentally-verifiable starting point for state manipulation. The single-particle operators
$\hat{c}_{\nu}$ $=$ $\sqrt{\frac{2}{L+1}}\sum_{k=1}^{L}\sin\left(q_{\nu}k\right)\hat{b}_{k},\phantom{a}q_{\nu}=\frac{\pi\nu}{L+1}$ (6)
diagonalize the Hamiltonian $\hat{H}=\sum_{\nu=1}^{L}\epsilon_{\nu}\hat{c}^{\dagger}_{\nu}\hat{c}_{\nu}$ with $\epsilon_{\nu}=-2J\cos\left(q_{\nu}\right)$. These operators satisfy the chirality relation $\hat{S}^{\dagger}\hat{c}_{\nu}\hat{S}=\hat{c}_{L-\nu+1}$ with $\epsilon_{\nu}=-\epsilon_{L-\nu+1}$.
From these $L$ single-particle states, we can construct a $d_{0}$-dimensional basis of $N$-particle non-interacting states with zero energy and definite chirality. First, note that for $L$ odd the single-particle state $|s\rangle\equiv c^{\dagger}_{(L+1)/2}|0\rangle$ has zero energy and chirality $\chi_{s}=-1$. Second, two-particle states of the form
$|p(\mu)\rangle\equiv\hat{S}^{\dagger}\hat{c}^{\dagger}_{\mu}\hat{S}\hat{c}^{\dagger}_{\mu}|0\rangle\ \mbox{for}\ \mu\in\lfloor L/2\rfloor$ (7)
have zero energy and chirality $\chi_{p}=+1$. There are precisely $d_{0}$ ways to distribute $N$ indistinguishable bosons among these chiral pair states $|p(\mu)\rangle$ and chiral single-particle states $|s\rangle$, and from these a standard basis for the zero energy subspace can be built.
Figure 2: Holonomies for adiabatic evolutions from $\theta=0$ to $2\pi$. (a) The holonomy matrix $\mathbf{G}$ in pair basis $|p(\mu)\rangle$ (Eq. (7)) reveals non-trivial adiabatic holonomies ( here $N=2$, $L=6$, and $N_{\theta}=10^{4}$ in Eq. (8), error $\sim\!10^{-3}$, see [62]). (b) The smallest integer $m_{\mathrm{diag}}$ such that $G^{m_{\mathrm{diag}}}$ is diagonal (within the error bounds) quantifies how nontrivial $\hat{G}$ is. $m_{\mathrm{diag}}$ increases strongly with system size. The two-body density of an initial zero energy state (c), here $|p(1)\rangle$, alters significantly during the adiabatic evolution (d).
Adiabatic manipulation of the zero energy subspace.— The periodic dependency in the statistical angle $\theta$ invites the question how the zero energy subspace is affected under cyclic adiabatic variation of $\theta$. As the statistical angle is tuned, the zero energy subspace $\mathcal{H}_{0}$ moves and rotates within the larger subspace $\mathcal{H}_{\pm}$ with majority chirality $\chi=\pm 1$. Similar to the Aharonov-Bohm effect on a ring, the zero energy subspace accumulates topological Berry phases and nontrivial holonomies. These act as unitary transformations on the above-given standard basis vectors when tuning $\theta$ from 00 to $2\pi$, as schematically depicted in Fig. 1(b3).
We propagate the zero energy subspace from $\theta=0$ to $\theta=2\pi$ using Kato’s adiabatic evolution [50, 63]
$\hat{G}=\lim_{N_{\theta}\to\infty}\prod_{j=1}^{N_{\theta}}\hat{P}_{j},$ (8)
where $\hat{P}_{j}=\sum_{\mu=1}^{d_{0}}|\psi_{\mu}(2\pi j/N_{\theta})\rangle\langle\psi_{\mu}(2\pi j/N_{\theta})|$ is the projector onto the zero energy subspace at the statistical angle $2\pi j/N_{\theta}$ and $|\psi_{\mu}\rangle$ is an orthonormal basis for the chiral zero energy subspace; see [62] for numerical details. The unitary matrix $\mathbf{G}_{\mu\mu^{\prime}}=\langle\psi_{\mu^{\prime}}|\hat{G}|\psi_{\mu}\rangle$ then embodies the adiabatic holonomy for a complete loop.
For $L=2$ the zero energy subspace is non-degenerate and we calculate the (Abelian) Berry phase $\mathbf{G}=\pm 1$ exactly and find the nontrivial value of $-1$ for consecutive odd-integer coefficients in the Gauss sum $\sum_{k=1}^{N/2}k$, see [62]. We investigate the non-Abelian holonomies of larger systems by exact diagonalization [62, 64, 65]. As an example, in Fig. 2(a), we present the holonomy matrix $\mathbf{G}$ in the chiral pair basis $|p(\mu)\rangle$ at $\theta=0$ for a system with $N=2$ and $L=6$ and $d_{0}=3$. The adiabatic evolution causes a significant non-trivial rotation of the zero energy subspace during the adiabatic evolution, which we indicate by ${\rho}_{ij}^{\nu}\left(\theta\right)=\langle\nu(\theta)|\hat{b}_{i}^{\dagger}\hat{b}_{j}^{\dagger}\hat{b}_{j}\hat{b}_{i}|\nu(\theta)\rangle$, an observable accessible in the corresponding experiments [31]. For example, we can take the initial two-body density $\rho_{ij}^{\nu}(\theta=0)$ of the basis state $|p(1)\rangle$ [Fig. 2(c)] and compare with ${\rho}_{ij}^{\nu}(\theta=2\pi)$, i.e., the two-body density corresponding to the adiabatically propagated state $\hat{G}|p(1)\rangle$. This yields a significant difference in the densities, see Fig. 2(d). Interestingly, for variations of parameters other than $\theta$ that preserve chiral symmetry, such as local variations in hopping strength $J$, we find that the connection is flat, meaning that $\hat{G}$ is not altered. Therefore, the results for the holonomy matrix $\mathbf{G}$ are robust against fluctuations in the adiabatic manipulation.
Tuning the statistical angle therefore implements a topologically protected operation on the zero energy subspace, a concept which has been extensively explored in the context of Majorana modes and other non-Abelian anyons [8]. To quantify the nontriviality of $\mathbf{G}$, we determine the smallest integer $m_{\mathrm{diag}}$ such that $\mathbf{G}^{m_{\mathrm{diag}}}$ is diagonal, i.e., trivial in the context of state manipulation. Also, at least $m_{\mathrm{diag}}$ states in the computational space can be prepared from an initial state by repeated application of this operation. For instance, the braiding of Majorana modes becomes diagonal for $m_{\text{Maj.}}=2$ and trivial with $G_{\text{Maj.}}^{4}=1$ [8]. Within our error bounds, we find $m_{\text{diag}}>4$ for various particle numbers and system sizes, particularly for large ones, see Fig. 2(b) and [62].
This idea can be elevated to non-Abelian holonomies if we implement the site-dependent statistical parameter $\theta_{j}$. This leads to a generalized anyonic exchange algebra $a^{\dagger}_{j}a_{k}=\hat{a}_{i}\hat{a}_{j}^{\dagger}-e^{i\theta_{i,j}}\hat{a}_{j}^{\dagger}\hat{a}_{i}=\delta_{i,j}$ with $\theta_{i,j}=\theta_{j}$ for $i>j$, $\theta_{i,j}=-\theta_{i}$ for $i<j$, and $\theta_{i,i}=0$. For different exchange phases $\theta_{L}$ in the left and $\theta_{R}$ in the right part of the system [62, 66], we generally find $\left[\mathbf{G}_{L},\mathbf{G}_{R}\right]\neq 0$, meaning the fluctuation-protected state manipulation depends on which statistical angle is altered first [62].
Figure 3: Steering to the zero energy subspace, starting with $N=2$ bosons at the center of $L=6$ sites. (a) The statistical phase $\theta(t)$ is dynamically varied along $60$ optimized interpolation points (gray) such that (b) the overlap $\mathcal{O}(t_{f})$ of the time-evolved state with the zero energy state $|p(1)\rangle$ (Eq. (7)) is maximal, reaching a final overlap of $1-10^{-4}$. The time evolution of (c) the chirality and (d) the static structure factor $C_{\pi}(t)$, which approaches its maximal value $N^{2}/L^{2}$ (dashed line) indicating the emergence of a checkerboard pattern, the signature of the zero energy subspace.
Steering to the zero energy subspace.— The basis states for $\theta=0$ can be prepared by temporal variation of $\theta=\theta(t)$ [31]. As a simple, accessible initial state, we consider two bosonic particles at the central sites of a lattice with $L=6$ sites, i.e., $|\Psi(t=0)\rangle=|001100\rangle$. In order to find a path for $\theta(t)$ which maximizes the overlap of the propagated wave function at time $t_{f}$ with a chiral pair target state, $\mathcal{O}(t_{f})=|\langle p(\mu)|\Psi(t_{f})\rangle|^{2}$, we represent $\theta(t)$ with $M$ interpolation points (for details see [62]). These $M$ points serve as an input for a gradient based optimization algorithm which updates the path $\theta(t)$ until $\mathcal{O}(t_{f})$ becomes maximal [67, 68, 69].
In Fig. 3, we present details of the optimization routine for $M\!=\!40$ and $t_{f}\!=\!40$. The chosen path for $\theta(t)$ shown in Fig. 3(a) reaches a fidelity $\mathcal{O}(t_{f})=1-10^{-4}$ in Fig. 3(b), which can be further increased for larger $M$ and $t_{f}$. As a measure whether the final state converges to the zero energy subspace, we monitor the expectation value of the chiral operator $\hat{S}$ in Fig. 3(c) and investigate the steering of the structure factor [70] in Fig. 3(d)
$C_{q}(t)=\frac{1}{L^{2}}\sum_{jk}e^{iq(j-k)}\langle\Psi(t)|\hat{n}_{j}\hat{n}_{k}|\Psi(t)\rangle.$ (9)
We find a dominant signal at $q=0,\pi$ that signals the presence of the checkerboard pattern in the $N$-particle density correlation for $L>N$, which is conveniently experimentally accessible in few-particle systems, see Fig. 2(c) [71, 72]. For large particle numbers, we instead propose to probe the one-body density matrix $\langle\psi|\hat{b}_{i}^{\dagger}\hat{b}_{j}|\psi\rangle$ directly by the Fourier transform of the quasi-momentum distribution $\langle\tilde{b}^{\dagger}_{k}\tilde{b}_{k}\rangle$.
Conclusions.— We have shown that chiral-symmetric number-preserving correlated hopping models with bosons accommodate at least $d_{0}$ degenerate zero energy states. The dimension of this space of states is robust against any parameter variations that preserve chiral symmetry, although the space itself sweeps through the chiral majority subspace. For the free anyon-Hubbard model, states in the zero energy chiral subspace can be prepared from experimentally-accessible initial states and detected by the characteristic checkerboard pattern in configuration space with current quantum gas microscopy techniques [31]. As the statistical angle is adiabatically tuned, the zero energy space picks up a non-Abelian holonomy for each cycle. This paradigm of non-Abelian state preparation could find further applications within the space of chiral symmetric correlated hopping models by tuning more than one cyclic parameter. As an example, we propose implementing this non-Abelian state preparation scheme within a spatially-inhomogeneous anyon-Hubbard model. More generally, we believe that Floquet-driven density-dependent Peierls phases offer a rich perspective for future exploration. Such models contain synthetic magnetic fluxes in few-body configuration space, and they provide an alternate path to understanding topological interactions in low dimensional systems. We propose a generalized anyon-Hubbard model with inhomogeneous statistical angles as one example of wider class of bosonic Hamiltonians with density-dependent Peierls phases, possibly without anyonic interpretation, but relevant to experiments and topological control via adiabatic holonomies.
The authors thank André Eckardt and Peter Græns Larsen for discussions on correlated hopping models and Joyce Kwan and Bryce Bakkani-Hassani for discussions of their anyon-Hubbard model experiment. T.P. thanks Ingo Runkel for comments on the local varying statistical angles. M.B. and T.P. acknowledge funding by the European Union (ERC, QUANTWIST, project number $101039098$). The views and opinions expressed are however those of authors only and do not necessarily reflect those of the European Union or the European Research Council, Executive Agency. T.P. and P.S. acknowledge support by the Cluster of Excellence “CUI: Advanced Imaging of Matter” of the Deutsche Forschungsgemeinschaft (DFG) – EXC 2056 – project ID $390715994$ and T.P. acknowledges funding of the DFG project No. 420120155. Additionally, N.H. was supported by the Deutscher Akademischer Austauschdienst, the Centre for Ultrafast Imaging: Advanced Imaging of Matter of the University of Hamburg.